News and Updates

News and Updates

About the blog

Read below our regular monthly / quarterly updates of work in progress and operations of the diesel locomotives based at the Bodmin and Wenford Railway.

Read more about the loco's and the group on our full website by clicking here.

May - June

UpdatesPosted by Administrator Fri, June 10, 2016 16:23:21

Loco Status

47306, P403D ‘Denise’ - In overhaul.

08444, D3452, 33110, 37142 & 50042 – In Traffic.

Operational dates this period.

14th May x5 return runs 37142 on the Red timetable.

22nd May x3 return runs 33110 on the Sunday Yellow TT.

Three days of mixed traction over the spring bank holiday weekend (28/29 & 30th May, 37142, 50042 & 33110 respectively). Though 50042 failed see below.

4th June x5 return runs 37142 (& 50042 for the last two on test) on the red timetable.

Various dates – 08444, D3452 & 33110 - Permanent way and station shunt duties.

Future running dates.

11th June x5 return runs (33110) Red timetable Diesel hauled all day.

28th July Evening Quiz night train, (33110 or 37142).

30th July x5 return runs (37142).

13th Aug x5 return runs (Loco TBC).

Maintenance for Service Duties


Exam work, found the B side battery box is deteriorating badly, which will soon require some fabrication work. Temporary remedial work to protect the batteries and to clean up / service the batteries has been undertaken until 33110 can be taken out of service for this work.

Gearbox on Bogie 2 topped up on new arrived gearbox ‘Motak’

Exams A, B & C continue as time allows.

33110 hauls away from Bodmin Parkway on the Bank Holiday Monday 30/5/16


Performed duties without fault.

50042 ‘Triumph’

In mid May ‘Triumph’ was extracted from Parkway shed. Serviced, run up and tested ready for it's mixed traffic turn and the following weekends EE Day. However during the Sunday mixed traffic duty, the Electric Train Heat Generator failed, this in turn meant that the Engine cooling radiator fan stopped working. 50042 had to be withdrawn early afternoon on a water temperature alarm and 37142 took over for the latter part of the day.

Bank Holiday Monday two volunteers gave up their day to try and diagnose the fault, which proved to be the KV10 (CU6, Not a small piece of 50042!), a stored spare (Of unknown condition) was dusted off and serviced, but it was too late to swap in on the Monday. A difficult decision was made knowing a large number of travellers were making plans to attend EE Day and no certainty that 50042 would be serviceable, meant we took the decision to postpone the event (Now cancelled).

More volunteers attended on the Thursday to carry out more tests and then swapped over the KV10 defective with the stored for 25 years out of service one. A lot of praying that components didn’t just fail after so long in storage worked and 50042 started generating ETHG again. Sat the 4th saw 37142 & 50042 out early on a test run, before 37142 took service. Some minor faults were worked through on 50042 during the morning and an A & B Exam undertaken, before the last two service trains saw a top & tail, followed by a double header were made.

Cab 1 DSD pedal stripped out and fixed from its sticking issue.

Cab 2 drivers wiper motor that failed again on the March DEC, fixed.

Cab 2 driver door has a new metal kicking strip.

Reverser was sticking, seals renewed and reverser tested.

4th June, 37142 & 50042 pass each other after service.

In Overhaul

47306 ‘The Sapper’

Work progress on 47306 with most weeks seeing a presence from the volunteers of two or three days out of each week even with the groups operational duties now peaking. Cab 1’s floor has also now been lifted the same as Cab 2 with preservation work being undertaken on the structural metal work underneath.

Boiler room work around the blower motor and tanks is nearly complete! With the blower motor back in position and preparations for a final paint out of this area with the reinstated pipework moves forward. The new boiler room roof panel has also been 1st fitted and now painted up.

With better outside working conditions, the task of stripping back and priming the buffer beams has got underway. Along with the removal of the buckled route marker plate from Cab 2 end (Not a quick task!), lamp units removed, it was decided a new plate was the best option. So one has been cut, drilled and now is being painted up ready for the refurbished route lamps to be mounted before being installed into the Cab bodywork.

Some preparation work has begun on the Cab 2 external body work around the roof and gutters. There has also been lots of painting of small parts of various hatch panels and other pieces of 47306.

  • Comments(0)//